# Measuring CEX-DEX Extracted Value and Searcher Profitability: The Darkest of the MEV Dark Forest

### v2

@@ 3.1 Current CEX-DEX Arbitrage Landscape
Despite the overall increase in CEX-DEX arbitrage activity, we observe reduced diversity of successful searchers over time. Before October 2024, 23 labeled searchers collectively accounted for nearly 99% of the total CEX-DEX arbitrage volume on-chain. This landscape transformed dramatically afterward, with just 14 labeled searchers remaining active and successful. The figure further reduced to 11 by Q1 2025, with three leading searchers capturing 90% of the total volume (cf. Appendix D Figure 15). Among all searchers, Wintermute and SCP have maintained dominant positions throughout the entire period. Interestingly, Kayle, initially a relatively minor searcher, began to rise significantly in influence around June 2024 and emerged as a leading searcher.
@@ 4.1 Searcher-Specific Optimal Execution Horizons
19 out of 23 labeled searchers follow the first two patterns and exhibit a characteristic optimal execution horizon, identified by their (median) gross return rising to or maintaining at its maximum before declining. These patterns align with theoretical expectations: before hedge execution begins, the spread typically increases as the trader’s information advantage manifests in favorable price movements. Once hedging commences, gross return declines as the trade’s market impact pushes prices against their position. For instance, we observe that the median gross return of SCP’s trades reaches its maximum $0.5$ seconds after the slot time, while the optimal horizon for Lucian and Maokai occurs at $1.0$ second, for Wintermute at $1.5$ seconds.
@@
We compute the profit margin only for trades whose estimated revenue is positive. Trades with $\widehat{\operatorname{EV}}_{i}\leq 0$ are labeled as ”N/A” for margin calculations and excluded from profit margin analysis, but they are retained in all other analyses related to PnL directly.
Our approach to inferring CEX-DEX searchers’ optimal execution horizon builds on methodologies from informed trading and market microstructure literature. Specifically, to identify this timing for each searcher, we adapt strategies from [12, 42], who empirically recover hidden execution timing of informed traders from observable post-trade price movements, by observing how CEX price movements affect the spread of their trades around the slot time. Moreover, our data-driven framework aligns with theoretical predictions by [7, 13], suggesting each trader possesses a unique, liquidity-driven optimal execution horizon characterized by a distinct ”peak-then-decay” pattern in returns (cf. Section 5.2). While our method necessarily provides a simplified approximation due to unobservable hedging price-impact costs, it nonetheless yields a consistent, empirically grounded upper-bound estimate of extracted value and PnL across searchers.
@@
However, for four searchers—Bard, Jinx, Tristana and Lux—their gross return curve is essentially flat (or drifts only marginally downward) across the $[-1,10]$ seconds window, so no clear ”peak-then-decay” point emerges. The most plausible interpretation is that these searchers hold their inventory and do not hedge their positions on CEX within this window: if they were unwinding inventory promptly, the resulting price impact on the CEX leg should depress the gross return. More explanations can be found in Section 5.2.
@@ 7.2 Searcher Profitability Analysis
Table 1 summarizes the profitability status of the searchers. Despite being market leaders, Wintermute, SCP, and Kayle clear the largest notional volume and earn the highest total revenues, yet their profit margin is among the lowest and their median PnL per trade mostly hovers only a few dollars above zero. SCP and Kayle’s total PnL are barely larger than that of mid-tier neutral searcher Shen, whose trade volume is much lower. Moreover, Graves, who sends almost all their trades to Titan, exhibits negative profitability according to our estimation, indicating payments to the builder often exceed their arbitrage revenue. We report the medians rather than the cumulative means for metrics of revenue per trade, PnL per trade, and profit margin to avoid the disproportionate influence of outlier trades that skew the means; the medians therefore represent the typical trade more faithfully.
@@ 7.3 Impact of Exclusive CEX-DEX Flow on Market Structures
Such exclusivity partnership largely contributes to the increasing centralization in the searcher market during this period and onward (cf. Appendix D, Figure 15), and had significant repercussions for competing builders, particularly rsync. As a result of Titan’s expansion, in Figure 10(b), we observe rsync’s market share fell below 5% by September 2024, effectively retreating from competition in the builder market. At the same time, Wintermute’s median profit margin quickly rebounded, indicating a strategic shift away from aggressively supporting rsync towards maximizing searcher profitability. Interestingly, beaverbuild’s market position and SCP’s profitability remained robust despite similar competitive pressures, likely buffered by the timely emergence of another exclusive searcher, Senna (See Appendix E for results not presented here).
@@ 8 Integrated Builder Profits Correction
We collect the Ultra Sound bid adjustments data [51] and MEV-Boost payloads data [31], and analyze builder profitability from block 17866488 to block 21998438 spanning the period from August 8, 2023, to March 8, 2025. Table 2 summarizes the profitability of beaverbuild, rsync, and Titan, considering integrated searcher profits (if any). “Builder Margin” presents the cumulative mean of profit margin if only builder profit for each block is considered, while “Aggregated Profit Margin” presents the cumulative mean of profit margin if considering both builder profit and searcher PnL for each block. Note that the “Total Searcher PnL” presented in this table only considers the searcher’s trades in the respective integrated builder’s blocks, excluding those in other blocks, to focus solely on their performance as builders. We detail the calculation method in Appendix G.
@@ 9 Discussion
Previous studies have discussed the impact of vertical integration, exclusive orderflow, and builder market centralization, highlighting concerns such as censorship and proposer loss [48, 63, 39]. Our analysis pushes the centralization narrative one step upstream: the CEX-DEX arbitrage layer is becoming increasingly centralized, with a limited number of searchers capturing most arbitrage opportunities and higher extracted value.
The two largest CEX-DEX searchers by volume and revenue—SCP and Wintermute—are each vertically integrated with leading block builders, beaverbuild and rsync, respectively. This vertical integration grants them a decisive strategic advantage in securing block inclusion for their arbitrage transactions. The third best searcher Kayle emerged as a significant market participant beginning around June 2024, likely following an exclusive arrangement with the second largest builder Titan. Collectively, these three searchers dominate approximately three-quarters of the total arbitrage volume and value extracted, reinforcing further centralization in both the builder and searcher markets.
@@
Recent research also highlights centralization concerns in DEX solver markets, where notably, Wintermute and SCP similarly dominate the solver markets on prominent platforms like CoWSwap and Uniswap X [64]. Taken together, these observations suggest that the entire MEV supply chain—from solver to searcher to builder—is becoming increasingly consolidated around a few dominant market participants.
@@
One proposed mitigation discussed in prior works [45, 44, 39, 36] involves shorter block times to reduce price movements between blocks and limit overall arbitrage opportunities. While this may lower the absolute value extractable from CEX-DEX arbitrages, it is less clear that shorter block times would meaningfully shift the market structure, given the sustained advantage held by dominant, vertically integrated players.
@@
Limitations of Our Approach.
Although our improved heuristics expand the coverage of identified CEX-DEX arbitrage transactions compared to prior methodologies [48, 39], certain arbitrages may remain undetected, resulting in potentially underestimating searcher revenues and PnL. Given the resources available, our methodology represents the best feasible solution at present. Further details are provided in Appendix F.
@@ Acknowledgments
The work is partially supported by The Latest in DeFi Research (TLDR) Fellowship 2024 funded by Uniswap Foundation. We appreciate Gaussian Process, Kevin Pang, Bill Zhang, Tom Zhao, Romain Butteaud for helpful discussions on the methodology. We thank Burak Öz and Max Resnick for insightful comments on this work. We also thank Alexander Tesfamichael from Ultra Sound Money and Žan Knafelc from Eden Network for their help with data collection.

### v3

@@ 3.1 Current CEX-DEX Arbitrage Landscape
Despite the overall increase in CEX-DEX arbitrage activity, we observe reduced diversity of successful searchers over time. Before October 2024, 23 labeled searchers collectively accounted for nearly 99% of the total CEX-DEX arbitrage volume on-chain. This landscape transformed substantially afterward, with just 14 labeled searchers remaining active and successful. The figure further reduced to 11 by Q1 2025, with three leading searchers capturing 90% of the total volume (cf. Appendix D Figure 15). Among all searchers, Wintermute and SCP have maintained dominant positions throughout the entire period. Interestingly, Kayle, initially a relatively minor searcher, began to rise significantly in influence around June 2024 and emerged as a leading searcher.
@@ 4.1 Searcher-Specific Optimal Execution Horizons
19 out of 23 labeled searchers follow the first two patterns and exhibit a characteristic optimal execution horizon, identified by their (median) gross return rising to or maintaining at its maximum before declining. These patterns align with theoretical expectations: before hedge execution begins, the spread typically increases as the trader’s information advantage manifests in favorable price movements. Once hedging commences, gross return declines as the trade’s market impact pushes prices against their position. For instance, we observe that the median gross return of SCP’s trades reaches its maximum 0.50.50.5 seconds after the slot time, while the optimal horizon for Lucian and Maokai occurs at 1.01.01.0 second, for Wintermute at 1.51.51.5 seconds.
@@
We compute the profit margin only for trades whose estimated revenue is positive. Trades with $\widehat{\operatorname{EV}}_{i}\leq 0$ are labeled as ”N/A” for margin calculations and excluded from profit margin analysis, but they are retained in all other analyses related to PnL directly. [^5] For example, consider a trade with an estimated revenue of -1 USD, for which the searcher pays 10 USD to the builder. This results in a net PnL of -11 USD. Although the trade incurs a loss, computing the profit margin would yield an illogical value of +1100%, which motivates its exclusion from margin-based analysis.
Our approach to inferring CEX-DEX searchers’ optimal execution horizon builds on methodologies from informed trading and market microstructure literature. Specifically, to identify this timing for each searcher, we adapt strategies from [13, 45], who empirically recover hidden execution timing of informed traders from observable post-trade price movements, by observing how CEX price movements affect the spread of their trades around the slot time. Moreover, our data-driven framework aligns with theoretical predictions by [8, 14], suggesting each trader possesses a unique, liquidity-driven optimal execution horizon characterized by a distinct “peak-then-decay” pattern in returns (cf. Section 5.2). While our method necessarily provides a simplified approximation due to unobservable hedging price-impact costs, it nonetheless yields a consistent, empirically grounded upper-bound estimate of extracted value and PnL across searchers.
@@
However, for four searchers—Bard, Jinx, Tristana and Lux—their gross return curve is essentially flat (or drifts only marginally downward) across the [1,10][-1,10][ - 1 , 10 ] seconds window, so no clear ”peak-then-decay” point emerges. The most plausible interpretation is that these searchers hold their inventory and do not hedge their positions on CEX within this window: if they were unwinding inventory promptly, the resulting price impact on the CEX leg should depress the gross return. More explanations can be found in Section 5.2.
@@ 7.2 Searcher Profitability Analysis
Table 1 summarizes the profitability status of the searchers. Despite being market leaders, Wintermute, SCP, and Kayle clear the largest notional volume and earn the highest total revenues, yet their profit margin is among the lowest and their median PnL per trade mostly hovers only a few dollars above zero. SCP and Kayle’s total PnL are barely larger than that of neutral searcher Shen, whose trade volume is much lower. Moreover, Graves, who sends almost all their trades to Titan, exhibits negative profitability according to our estimation, indicating payments to the builder often exceed their arbitrage revenue. We report the medians rather than the cumulative means for metrics of revenue per trade, PnL per trade, and profit margin to avoid the disproportionate influence of outlier trades that skew the means; the medians therefore represent the typical trade more faithfully.
@@ 7.3 Impact of Exclusive CEX-DEX Flow on Market Structures
This dynamic played a major role in accelerating centralization within the searcher market during this period and beyond (cf. Appendix D, Figure 15), and had significant repercussions for competing builders, particularly rsync. As a result of Titan’s expansion, in Figure 10(b), we observe rsync’s market share fell below 5% by September 2024, effectively retreating from competition in the builder market. At the same time, Wintermute’s median profit margin quickly rebounded, indicating a strategic shift away from aggressively supporting rsync towards maximizing searcher profitability. Interestingly, beaverbuild’s market position and SCP’s profitability remained robust despite similar competitive pressures, likely buffered by the timely emergence of another exclusive searcher, Senna (See Appendix E for results not presented here).
@@ 8 Integrated Builder Profits Correction
We collect the Ultra Sound bid adjustments data [54] and MEV-Boost payloads data [33], and analyze builder profitability from block 17866488 to block 21998438 spanning the period from August 8, 2023, to March 8, 2025. Table 2 summarizes the profitability of beaverbuild, rsync, and Titan, considering integrated searcher profits (if any). “Builder Margin” presents the cumulative mean of profit margin if only builder profit for each block is considered, while “Aggregated Profit Margin” presents the cumulative mean of profit margin if considering both builder profit and searcher PnL for each block. Note that the “Total Searcher PnL” presented in this table only considers the searcher’s trades in the respective integrated builder’s blocks, excluding those in other blocks, to focus solely on their performance as builders. Recall, given that the searcher PnL is an upper-bound estimtate, the aggretated profit and margin are also upper-bound estimates. We detail the calculation method in Appendix G.
@@ 9 Discussion
Limitations of Our Approach.
Although our heuristics expand the coverage of identified CEX-DEX arbitrage trades compared to prior methodologies [51, 41], certain trades may remain undetected, resulting in potentially underestimating searcher revenues and PnL. Given the resources available, our methodology represents the best feasible solution at present.
Centralization in the searcher market.
Previous studies have discussed the impact of vertical integration, exclusive orderflow, and builder market centralization, highlighting concerns such as censorship and proposer loss [51, 67, 41]. Our analysis pushes the centralization narrative one step upstream. Critically, we are the first to provide robust quantitative estimates of the value extracted through CEX-DEX arbitrages and shed empirical light on profit-sharing dynamics previously understood only qualitatively.
It appears that, just as arbitrages in traditional finance have been concentrated among a few specialist firms [55]—often because they purchase retail order flow (i.e., payment for order flow) or control clearing licenses [26, 44]—arbitrages at the TradFi and DeFi interface are similarly centralized by searcher-builder orderflow arrangements and vertical integration.
Specifically, the two largest CEX-DEX searchers by volume and revenue—SCP and Wintermute
are each vertically integrated with leading block builders, beaverbuild and rsync, respectively. This vertical integration grants them a decisive strategic advantage in securing block inclusion for their arbitrage transactions. The third best searcher Kayle emerged as a significant market participant beginning around June 2024, likely following an exclusive arrangement with the second largest builder Titan. Collectively, these three searchers dominate approximately three-quarters of the total arbitrage volume and value extracted, reinforcing further centralization in both the builder and searcher markets.
@@
Recent research also highlights centralization concerns in DEX solver markets, where notably, Wintermute and SCP similarly dominate the solver markets on prominent platforms like CoWSwap, Uniswap X, and 1inchFusion [68]. Taken together, these observations suggest that the entire MEV supply chain—from solver to searcher to builder—is becoming increasingly consolidated around a few dominant market participants.
@@
One proposed mitigation discussed in prior works [48, 47, 41, 38] involves shorter block times (e.g., EIP-7782 [7]) to reduce price movements between blocks and limit overall arbitrage opportunities. While this may lower the absolute value extractable from CEX-DEX arbitrages, it is less clear that shorter block times would meaningfully shift the market structure, given the sustained advantage held by dominant, vertically integrated players.
@@ Acknowledgments
The work is partially supported by The Latest in DeFi Research (TLDR) Fellowship 2024 funded by Uniswap Foundation. We appreciate Gaussian Process, Kevin Pang, Bill Zhang, Tom Zhao, and Romain Butteaud for helpful discussions on the methodology. We thank Burak Öz and Max Resnick for insightful comments on this work. We also thank Alexander Tesfamichael from Ultra Sound Money and Žan Knafelc from Eden Network for their help with data collection.